### v1

# Peer-Aware Cost Estimation in Nonlinear General-Sum Dynamic Games for Mutual Learning and Intent Inference

Abstract
Human-robot interactions can be modeled as incomplete-information general-sum dynamic games since the objective functions of both agents are not explicitly known to each other. However, solving for equilibrium policies for such games presents a major challenge, especially if the games involve nonlinear underlying dynamics. To simplify the problem, existing work often assumes that one agent is an expert with complete information about its peer, which can lead to biased estimates and failures in coordination. To address this challenge, we propose a nonlinear peer-aware cost estimation (N-PACE) algorithm for general-sum dynamic games. In N-PACE, using iterative linear quadratic (LQ) approximation of the nonlinear general-sum game, each agent explicitly models the learning dynamics of its peer agent while inferring their objective functions, leading to unbiased fast learning in inferring the unknown objective function of the peer agent, which is critical for task completion and safety assurance. Additionally, we demonstrate how N-PACE enables intent communication in such multi-agent systems by explicitly modeling the peer’s learning dynamics.
I INTRODUCTION
General-sum dynamic games offer a powerful framework for modeling multi-agent interactions [1, 2], with applications in autonomous driving [3], assistive robots [4], and social navigation [5]. However, the process of finding the feedback Nash equilibrium (FBNE) solution of such games, which can be achieved by solving a set of Hamilton-Jacobi-Bellman (HJB) equations [6], is inherently complex, as traditional dynamic programming approaches suffer from the curse of dimensionality [7]. This problem becomes even more challenging when agents do not know each other’s objective function and must perform inference online while making decisions, resulting in an incomplete-information differential game, a common scenario in human-robot interaction (HRI) [4, 8, 9, 10] and autonomous vehicle interaction [11, 12, 13].
Figure 1: Demonstration of the core idea in N-PACE using a simulated general-sum game between two autonomous cars at an intersection. In (a), each agent assumes its peer knows its aggressiveness level, leading to biased inferences and possible collision. In (b), N-PACE allows agents to learn their peer’s aggressiveness while accounting for mutual learning, ensuring safe intersection crossing.
Existing approaches to solve such games often rely on the assumption that one agent is an expert which knows the true intent of other agents [14, 15, 1]. Despite some empirical success of these approaches, in scenarios where all agents are learning simultaneously and inferring each other’s objective functions (which is defined as their intent in this paper), these assumptions can lead to biased estimates, poor coordination, and failure to achieve each agent’s goal [16]. This is because while doing the inference, they fail to account for the fact that other agents can be non-stationary learners, whose policies are conditioned on previously observed actions [9]. It is even more important to consider such mutual learning in a competitive setting where an agent has no information of the others’ true intent.
In this paper, we highlight the importance of modeling the learning dynamics of peer agents while inferring their objectives, addressing critical limitations in existing approaches. We achieve this goal by developing a peer-aware cost estimation algorithm for general-sum nonlinear games (N-PACE), where each agent leverages prior knowledge about its peer agent’s learning dynamics (e.g., gradient-based learner and Bayesian learner) to condition its current inference of the other’s objective on their parameter learning, thereby removing the bias in their inferences. The N-PACE algorithm employs iterative linear quadratic (iLQ) approximation of the nonlinear general-sum game [5], both in policy update and intent inference phases. We also demonstrate how (iLQ) enables efficient approximation during the inference phase of our algorithm.
Contributions. This paper makes the following contributions. (1) We introduce the N-PACE algorithm for mutual intent inference in multi-agent interaction where all agents are making inference online. (2) We empirically demonstrate that N-PACE is critical for task completion and safety in multi-agent systems where all agents are learners, compared to the incorrect assumption that a peer agent is an expert who knows the ego agent’s true intent. (3) We demonstrate how N-PACE, along with access to the learning dynamics of the peer agent, can be leveraged for intent communication and signaling to facilitate mutual learning and adaptation.
The closest works to ours are [5, 15, 16], where [5] uses iLQ for solving complete information multi-agent interactions. In [15], the iLQ approach is employed in a general-sum game where an “expert” demonstrates its intent to a learner agent knowing their learning dynamics, but it does not consider the situation when two learning agents (no experts) try to coordinate which is a more practical setting in multi-agent interactions. Finally, while [16] studies the problem of mutual learning and adaptation, its problem setup is limited to a single-step optimization in the cost function. This work extends it to a finite-horizon optimization problem which is more common in multi-agent interactions.
II Related Work
Solving general-sum games. Multi-agent interactions have been explored using various frameworks, including multi-agent reinforcement learning [17, 18], adaptive control [19], and game theory [20, 21]. A large portion of these interactions can be modeled as general-sum dynamic games [22, 23]. Traditional approaches for solving such games rely on adaptive dynamic programming to compute feedback Nash equilibria (FBNE) [24]. However, these methods are often hindered by the curse of dimensionality [7], making them impractical for high-dimensional problems and real-time implementation in robotics. To overcome these limitations, physics informed neural networks (PINN) have been used to solve the (HJB) equation [23] more efficiently. Some recent work has focused on iterative linear-quadratic (iLQ) approximations, which offer a more computationally efficient strategy for for approximating the multi-agent interactions [5] solution by finding the local Nash equilibrium of such games. In this paper, we employ iLQR for both the learning and control phases to find the approximated Nash equilibrium policies.
Incomplete information general-sum games. Solving incomplete information general-sum games is significantly more challenging than their complete-information counterparts [25, 26]. These games often require simplifications, such as modeling one agent as a learner while assuming others are fully informed [14, 27], effectively simplifying them as optimal or adaptive control problems [4]. Other approaches discretize the intent parameter and rely on offline datasets for intent inference and belief updates [28, 3]. Our work is motivated by the lack of studies on incomplete information dynamic games where all agents are learners [29]. We highlight the potential issues caused by these simplifications, such as assuming other agents are fully informed.
Learning from a learner. Online inference of an expert agent’s objective function has been extensively studied through inverse optimal control and inverse reinforcement learning [30, 31, 32]. While inverse non-cooperative games have been explored both offline and online [33, 34], they also primarily focus on learning from an expert system. Some works have considered learning from a learner agent by incorporating their learning dynamics [35, 29], but they do not directly address interactions where multiple learner agents simultaneously infer each other’s cost or reward function parameters. Building on these learning from a learner approaches and recent efforts in modeling human learning dynamics for improved intent inference in human-robot interactions [10, 9], we propose a framework for mutual intent inference in systems where all agents are learners.
III Methods
In this section, we first introduce the notations and formulate the multi-agent interaction problem as a general-sum incomplete information differential game. We then discuss how iLQR can be used for solving such games. Next, we propose a framework for modeling each agent’s learning dynamics. Finally, we present our algorithm, N-PACE, and demonstrate its role in intent communication.
III-A Problem Formulation
Notations and Assumptions. Without loss of generality to more agents, we focus on a two-player setting with agents $i$ and $j$. When referring to an arbitrary agent, we use the index $k$ and denote their counterpart as $-k$. We consider the following system dynamics
$s_{t+1}=f_{t}(s_{t},a^{i}_{t},a^{j}_{t}),$ (1)
where $s_{t}\in\mathbb{R}^{n}$ is the shared state vector at time step $t$, and $a^{i}_{t},a^{j}_{t}\in\mathbb{R}^{m}$ are the control inputs of agents $i$ and $j$, respectively. The function $f_{t}$ is assumed to be differentiable with respect to $s_{t},a^{i}_{t},$ and $a^{j}_{t}$. We assume that all states and control signals $s_{t},a^{i}_{t},a^{j}_{t}$ are observable to both players at each time step. For a fixed time horizon $T$, the running cost of each player $k\in\{i,j\}$ is defined as $g^{k}_{t}(s_{t},a^{i}_{t},a^{j}_{t};\theta^{k})$, where $g^{k}_{t}$ is a scalar function that is twice differentiable with respect to $s_{t},a^{i}_{t},a^{j}_{t}$ and differentiable with respect to the intent parameter $\theta^{k}$ for $g^{k}_{t}$ (e.g., the targeted goal state or a weighting factor). Additionally, we use the notation $\langle\hat{\cdot}\rangle$ to denote the parameter estimate.
General-sum game formulation. We consider a receding horizon cost minimization problem for each agent, where at each time step $t\in\{0,1,\dots,T\}$, each player $k$ aims to minimize the cumulative cost:
$J^{k}_{t}(s_{t},a^{i}_{t},a^{j}_{t})=\sum_{\tau=t}^{T}g^{k}_{\tau}(s_{\tau},a^{i}_{\tau},a^{j}_{\tau};\theta^{k}).$ (2)
Feedback Nash Equilibrium (FBNE) [36] is a solution concept for the above minimization problem, constrained by (1), where neither agent has an incentive to deviate from their policy, as doing so would increase their cumulative cost. In an FBNE, each agent’s policy is a function of the current state, ensuring that at every time step, the chosen control actions are optimal given the strategies of the other agents. We represent these policies using $\pi^{k}_{t}(s_{t};\theta^{i},\theta^{j})$ such that $a^{k}_{t}=\pi^{k}_{t}(s_{t};\theta^{i},\theta^{j})$, emphasizing that they depend on both $\theta^{i}$ and $\theta^{j}$, which is crucial for further formulation in our work.
The incomplete information game. The previously discussed FBNE solution was derived under the assumption that both agents have full knowledge of $\theta^{i}$ and $\theta^{j}$. However, in incomplete information games, $\theta^{j}$ may be unknown to agent $i$, and $\theta^{i}$ may be unknown to agent $j$. As a result, at each time step, they must rely on estimations $\hat{\theta}^{j}_{t}$ and $\hat{\theta}^{i}_{t}$, respectively, and update these estimates over time. This leads to policies of the form $\pi^{i}_{t}(s_{t};\theta^{i},\hat{\theta}^{j}_{t})$ and $\pi^{j}_{t}(s_{t};\hat{\theta}^{i}_{t},\theta^{j})$, which are not necessarily the same as FBNE strategies. To address this problem, the next section introduces how we use the iLQR method as a solver to obtain the policy $\pi^{k}_{t}(s_{t};\theta^{k},\hat{\theta}^{-k}_{t})$ for each agent $k$, based on their current estimate of the other agent’s parameter $\hat{\theta}^{-k}_{t}$. We then propose appropriate update rules for these estimates in Sec. III-C.
III-B iLQR for Multi-Agent Interactions
The iterative linear quadratic regulator (iLQR) approach for locally optimal feedback control of nonlinear systems was first introduced in [37]. More recent work, such as [5], has demonstrated its efficiency in computing local Nash equilibrium solutions for general-sum dynamic games in real time. In our incomplete information game, where we need policies $\pi^{k}_{t}$ to be computed as functions of the estimated parameters $\hat{\theta}^{i}_{t}$ and $\hat{\theta}^{j}_{t}$ in real-time, we leverage the proposed iLQR algorithm in [5] to solve the game for each agent $k$ at every time step. The main reasons for using iLQR in this work are its efficiency in real-time computations and, more importantly, its capability to parameterize the obtained policies as functions of $\hat{\theta}^{i}_{t}$ and $\hat{\theta}^{j}_{t}$, which is crucial for N-PACE as shown in Sec (III-D).
In the iLQR approach, we begin with an initial admissible policy for each agent. In the forward pass, by applying these policies to the system dynamics in (1), we obtain a trajectory $\boldsymbol{\zeta^{0}_{t}}=\{(s^{0}_{\tau},a^{i0}_{\tau},a^{j0}_{\tau})\mid\tau\in\{t,t+1,\dots,T\}\}$. Next, by linearizing the dynamics $f_{t}$ and quadratizing the running costs $g^{k}_{t}$ in the neighborhood of the trajectory $\boldsymbol{\zeta^{0}_{t}}$, we construct a finite-horizon LQ game, which can be solved via coupled Riccati equations using the methods proposed in [36]. The linear feedback control solution of this LQ game is then used to update the agents’ policies, as suggested in [5]. This iterative process continues until the system trajectory remains unchanged at each iteration and converges to a fixed trajectory, known to be the local Nash equilibrium trajectory $\boldsymbol{\zeta^{*}_{t}}$.
In an incomplete information game, both agents are learning the other agent’s parameter $\theta^{-k}$, resulting in an estimate $\hat{\theta}^{-k}_{t}$. As a result, each agent independently solves a different iLQR problem to update their policy using $\theta^{k}$ and $\hat{\theta}^{-k}_{t}$, leading to a distinct converged trajectory $\zeta^{*k}_{t}(\theta^{k},\hat{\theta}^{-k}_{t})$ at each step. At the converged trajectories, each agent’s policy consists of a feedforward time-varying term $\alpha_{t}$ and a feedback time varying term $P_{t}$ in the form of
$\pi^{k}_{t}(s_{t};\theta^{k},\hat{\theta}^{-k}_{t})$ $=-\alpha_{t}(\theta^{k},\hat{\theta}^{-k}_{t})$ (3)
$\quad-P_{t}(\theta^{k},\hat{\theta}^{-k}_{t})(s_{t}-s^{*k}_{t}(\theta^{k},\hat{\theta}^{-k}_{t})),$
where $s^{*k}_{t}(\theta^{k},\hat{\theta}^{-k}_{t})$ represents states stored in the converged trajectory. In (3), we explicitly maintain and emphasize the dependence of each parameter on $\theta^{k}$ and $\hat{\theta}^{-k}_{t}$. This not only highlights how agents update their policies based on new estimates but also serves as an introduction to the N-PACE algorithm.
Remark 1. The proposed framework in this paper is not limited to using iLQR for solving general-sum games. The key requirement for the success of N-PACE is to have a policy explicitly parameterized by $\theta^{k}$ and $\theta^{-k}$. In fact, alternative approaches, such as some reinforcement learning algorithms [38, 17], can also be used.
III-C Modeling the Learning Dynamics
In Sec. III-B, we discussed how an agent $k$ can update its policy at each step based on the current estimate $\hat{\theta}^{-k}_{t}$ by solving a general-sum game. However, the question of how to learn and update $\hat{\theta}^{-k}_{t}$ remains open. In this section, we propose our learning algorithm under the assumption that agent $-k$ is an expert, for the sake of clarity and brevity. We then introduce approximation methods to enhance the efficiency of this learning process for real-time applications. Finally, in Sec. III-D, we discuss how this learning approach can be extended to N-PACE for scenarios where all agents are learners.
The learning dynamics $h^{k}_{l}$. When agents can observe the states and control signals $s_{t},a^{k}_{t},a^{-k}_{t}$ at each step, they can leverage this information to construct an error term for the online learning of the parameters $\hat{\theta}^{-k}_{t}$. An agent $k$, knowing that their peer is solving an iLQR game to generate policies, can attempt to model this policy using their current estimate $\hat{\theta}^{-k}_{t}$ as $\hat{a}^{-k}_{t}=\pi^{-k}_{t}(s_{t};\hat{\theta}^{-k}_{t},\theta^{k})$ as given in (3). Consequently, they can define the following update rule as their learning dynamics $h^{k}_{l}$ to refine the estimate of $\hat{\theta}^{-k}_{t}$
$\hat{\theta}^{-k}_{t+1}$ $=h^{k}_{l}(\hat{\theta}^{-k}_{t},\theta^{k},a^{-k}_{t},s_{t})=\hat{\theta}^{-k}_{t}$ (4)
$-\alpha_{k}(\pi^{-k}_{t}(s_{t};\hat{\theta}^{-k}_{t},\theta^{k})-a^{-k}_{t})^{\top}\nabla_{\hat{\theta}^{-k}_{t}}\pi^{-k}_{t}(s_{t};\hat{\theta}^{-k}_{t},\theta^{k})$
Remark 2. Notice that in (4), when modeling the policy of the other agent $-k$, the parameter $\theta^{k}$ is included in the policy model. This effectively assumes that the other agent has full knowledge of $\theta^{k}$ and is treated as an expert, which can lead to incorrect and biased estimates as shown in the case studies of this paper.
Remark 3. The proposed learning dynamics $h^{k}_{l}$ follows a simple gradient descent algorithm with a learning rate $\alpha_{k}$, but it is not restricted to gradient-based methods in the N-PACE. It can also take the form of a Bayesian learner or even a neural network model, as explored in [10]. In fact, we implement a Bayesian inference update in all our case studies in Sec. IV, assuming a Gaussian observation model for parameter estimation, to achieve more robust estimation results.
Efficient approximation of the gradient. The learning dynamics proposed in (4) require calculating the gradient of the policy model with respect to the unknown parameter $\hat{\theta}^{-k}_{t}$. The iLQR method, which generates the policy at each step, is an iterative process where the generated trajectories and policy improvement steps depend on the unknown parameter $\hat{\theta}^{-k}_{t}$. While this approach still relates the policy to $\hat{\theta}^{-k}_{t}$, the iterative nature of iLQR can make gradient computation computationally expensive or even intractable for real-time implementation. Notably, in (3), as well as in the converged trajectory $\zeta^{*k}_{t}(\theta^{k},\hat{\theta}^{-k}_{t})$, all parameters are explicitly expressed as functions of $\hat{\theta}^{-k}_{t}$.
To address this challenge, we adopt the approximation method proposed in [34]. Specifically, we ignore the dependency of the converged trajectory in iLQR on $\hat{\theta}^{-k}_{t}$ and replace the gradient of the policy with the gradient of the linear policy obtained by solving the LQ game around the converged trajectory of iLQR. The intuition behind this approximation is that in iLQR, the policy derived from solving the expanded LQ game around the converged trajectory serves as a good approximation of the final converged policy itself, as discussed and demonstrated in [34]. However, automatic differentiation (AD) methods can still be used to compute this gradient if there are no computational constraints. Alternatively, if a different game solver allows the policy to be directly expressed as a function of the unknown parameter $\hat{\theta}^{-k}_{t}$ (e.g., an actor network receiving the parameter as a state input), one can bypass the need for this approximation method.
III-D N-PACE for Multi-Agent Interactions
In Sec. III-C, we proposed the learning dynamics $h^{k}_{l}$ for each agent. However, the update rule in (4) assumes that agent $-k$ is aware of $\theta_{k}$ while generating its policy, effectively treating agent $-k$ as an expert during estimation. This assumption is clearly invalid when all agents are learning. In N-PACE, we address this issue by ensuring that each agent $k$ not only learns its own parameters but also accounts for the learning procedure of the other agent $-k$. This is feasible when agents have access to both learning dynamics, $h^{i}_{l}$ and $h^{j}_{l}$. Consequently, they can replace $\theta_{k}$ with $\hat{\theta}^{k}_{t}$ in equation (4), eliminating the bias in the learning process. As a result, each agent, by having access to learning dynamics and tracking the learning of the other agent, arrives at the following system of learning equations at each step
$\hat{\theta}^{i}_{t+1}$ $=h^{j}_{l}(\hat{\theta}^{i}_{t},\hat{\theta}^{j}_{t},a^{i}_{t},s_{t}),$ (5)
$\hat{\theta}^{j}_{t+1}$ $=h^{j}_{l}(\hat{\theta}^{j}_{t},\hat{\theta}^{i}_{t},a^{j}_{t},s_{t}).$
An important property of N-PACE is that when both agents start with the same initial estimates for $\hat{\theta}^{i}_{0}$ and $\hat{\theta}^{j}_{0}$ and continuously track each other’s learning, they will always track the same estimations for both parameters. As a result, the learning dynamics (5) is effectively centralized, even though each agent performs the learning process independently. The assumption of agreeing on initial estimates is not restrictive in many multi-agent interaction scenarios. This holds true when there is a strong prior on the parameters (e.g., a predefined level of rationality or aggressiveness in autonomous driving) or when agents are designed to coordinate despite having different and initially unknown intentions. Even if the initial estimates or learning dynamics of each agent are uncertain (such as human-robot interactions), we refer the readers to works such as [10], which focus on modeling these learning processes and initial estimates in real time, or offline data-driven approaches for modeling the learning dynamics. Alg. 1 summarizes all the steps needed for N-PACE performed by agent $i$.
Algorithm 1 N-PACE for agent $i$
Initialization: Initialize $\hat{\theta}^{j}_{0}$ and $\hat{\theta}^{i}_{0}$.
for each time step $t<T$ do
     1. Policy Generation:
       a. Using (3), generate the policy as:
       $\pi^{i}_{t}(s_{t};\theta^{i},\hat{\theta}^{j}_{t})=iLQSolver(s_{t},\theta^{i},\hat{\theta}^{j}_{t})$
       b. Apply the control $a^{i}_{t}=\pi^{i}_{t}(s_{t};\theta^{i},\hat{\theta}^{j}_{t})$
     2. Prediction:
       a. Predict the $j$’s policy as $\pi^{j}_{t}(s_{t};\hat{\theta}^{j}_{t},\hat{\theta}^{i}_{t})$ and model
       how $j$’s predict your policy $\pi^{i}_{t}(s_{t};\hat{\theta}^{i}_{t},\hat{\theta}^{j}_{t})$ using the
       centralized solver $iLQSolver(s_{t},\hat{\theta}^{i}_{t},\hat{\theta}^{j}_{t})$
       b. compute the gradient of each predicted policy
       with respect to their corresponding parameter $\nabla_{\hat{\theta}^{k}_{t}}{\pi^{k}_{t}}$
     3. Learning:
       a. observe $s_{t},a^{j}_{t}$, form the prediction errors.
       b. Using the prediction and computed gradients apply
       the learning dynamics as: $\hat{\theta}^{i}_{t+1}=h^{j}_{l}(\hat{\theta}^{i}_{t},\hat{\theta}^{j}_{t},a^{i}_{t},s_{t})$
       $\hat{\theta}^{j}_{t+1}=h^{j}_{l}(\hat{\theta}^{j}_{t},\hat{\theta}^{i}_{t},a^{j}_{t},s_{t})$
end for
III-E N-PACE for Intent Communication
The N-PACE algorithm also enables intent communication. Notice that the learning dynamics of agent $-k$, denoted as $h^{-k}_{l}$, are fundamentally conditioned on the actions $a^{k}_{t}$ taken by agent $k$. As a result, each agent can modify its actions to influence the learning process of the other. To enable this, we propose to introduce an additional intent communication term in the form of $\eta(\hat{\theta}^{k}_{t}-\theta^{k})^{2}$ to the instantaneous cost function of each agent, $g^{k}_{\tau}(s_{\tau},a^{i}_{\tau},a^{j}_{\tau};\theta^{k})$, where $\eta$ is the weighting scalar. This leads to the following optimization problem for each agent $k$:
$\min_{a^{k}_{t}}$ $J^{*k}_{t}(s_{t},a^{k}_{t},a^{-k}_{t})=\sum_{\tau=t}^{T}g^{k}_{\tau}(s_{\tau},a^{k}_{\tau},a^{-k}_{\tau};\theta^{k})+\eta\|\hat{\theta}^{k}_{\tau}-\theta^{k}\|^{2}$ (6)
s.t. $s_{\tau+1}=f_{t}(s_{\tau},a^{k}_{\tau},a^{-k}_{\tau}),\ \hat{\theta}^{-k}_{{\tau}+1}=h^{k}_{l}(\hat{\theta}^{-k}_{\tau},\hat{\theta}^{k}_{\tau},a^{-k}_{\tau},s_{\tau}),$
$\hat{\theta}^{k}_{{\tau}+1}=h^{-k}_{l}(\hat{\theta}^{k}_{\tau},\hat{\theta}^{-k}_{\tau},a^{k}_{\tau},s_{\tau}).$
The influence problem above is computationally expensive and may be intractable for real-time implementation. Remember that the learning dynamics $h^{k}_{l},h^{-k}_{l}$ themselves require solving a general-sum game, using iLQR in this case, for their prediction phase and computing the gradient of that prediction. To address this challenge, we propose an approximation method for solving the optimization problem in (6). Instead of considering the evolution of the other agent’s parameter learning throughout the entire horizon $[t,T]$ in (6) to form the cumulative cost $J^{*k}$, we focus only on a single-step update of the other agent’s learning as $\hat{\theta}^{-k}_{{t}+1}=h^{k}_{l}(\hat{\theta}^{-k}_{t},\hat{\theta}^{k}_{t},a^{-k}_{t},s_{t})$, so we can modify the cumulative cost of agent $k$ at each step as
$J^{*k}(s_{t},\hat{\theta}^{-k}_{t},a^{i}_{t},a^{j}_{t})=J^{k}(s_{t},a^{i}_{t},a^{j}_{t})+\eta\|\hat{\theta}^{k}_{t+1}-\theta^{k}\|^{2},$ (7)
where $\hat{\theta}^{k}_{t+1}$ depends on $a^{k}_{t}$ through the learning dynamics $h^{-k}_{l}$. We will show the effectiveness of this approximation for intent communication in Case Study 3 in Sec. IV.
IV Case Studies
In this section, we demonstrate the importance of modeling the learning dynamics of the other agent using the proposed N-PACE for task completion, safety, and signaling in multi-agent systems. We illustrate this through three simulation case studies: 1) an assistive nonlinear lunar lander, 2) an autonomous car in lane merging, and 3) an autonomous car in intersection crossing. The learning dynamics in all case studies employ gradient descent-based Bayesian inference using a Gaussian belief.
IV-A Case Study 1: Assistive Lunar Lander
In this example, we study a shared control problem to show the importance of N-PACE in task completion. An assistive lunar lander scenario in which a human agent controls only the angle of the lunar lander $\phi$ by applying the torque $T$, while an autonomous agent controls the thrust force $F$ along the direction of the lunar lander’s angle (assuming no gravity). Both agents cooperate to land the lunar lander at the final position $(x_{f},y_{f})$, but with asymmetric information: the human knows only the final desired $x$-position $x_{f}$, while the autonomous agent knows only the final desired $y$-position $y_{f}$. As a result, the human must learn the parameter $\hat{y}_{f}$, and the autonomous agent must learn $\hat{x}_{f}$ based on observed interactions. The underlying nonlinear dynamics and the running costs for the human $g_{h}$ and the agent $g_{a}$ are defined as:
$\dot{x}$ $=v_{x},\quad\dot{y}=v_{y},\quad\dot{\phi}=\omega,\quad\dot{v}_{x}=F\sin{\phi},\quad\dot{v}_{y}=F\cos{\phi},$
$\dot{\omega}$ $=T,\quad g_{a}=10(x-\hat{x}_{f})^{2}+10(y-y_{f})^{2}+10(\phi-\frac{\pi}{2})^{2}$
$+10(v_{x}^{2}+v_{y}^{2}+\omega^{2})+F^{2},\quad g_{h}=10(x-x_{f})^{2}+$
$10(y-\hat{y}_{f})^{2}+10(\phi-\frac{\pi}{2})^{2}+10(v_{x}^{2}+v_{y}^{2}+\omega^{2})+T^{2}.$
The game has a time horizon of $T=5s$ with a sampling time of $\Delta T=0.1s$ that is used to discretize the above dynamics. In this example, we set $x_{f}=5$ and $y_{f}=-5$, while the lunar lander starts at the initial position $(-5,8)$ with a zero initial angle. We conducted three simulations for this case study: 1) a complete information game where both agents have full knowledge of $x_{f}$ and $y_{f}$ as ground truth, 2) a scenario where each agent treats its peer as an expert, assuming the other agent has complete information and 3) a scenario in which agents operate according to N-PACE.
Figure 2 shows summarizes the results, where Fig. 2 (a) presents the trajectory of the lunar lander under different algorithms, explicitly demonstrating the superiority of N-PACE compared to the approach where each agent treats its peer as an expert. It can be seen that the final position of the lunar lander in N-PACE is closer to the desired final position, and the entire trajectory aligns more closely with the complete information game trajectory (the local Nash equilibrium). In contrast, treating the other agent as an expert results in the lunar lander landing at a different position, underscoring the importance of modeling the learning dynamics of the other agent for successful task completion. Figs. 2 (b) and (c) present the results of parameter learning performance for each agent under different algorithms, illustrating how N-PACE successfully converges to the true final position values for each agent, whereas treating the other agent as an expert results in biased parameter learning. Notably, treating the other agent as an expert can lead to a significant overshoot in the learning process. Intuitively, this may be due to a high initial bias in assuming that the peer agent has complete information.
Figure 2: Results of the assistive lunar lander example. (a) Trajectory comparison between the complete information game, N-PACE, and the expert peer assumption. (b) and (c) show each agent’s learning performance for the final desired position under different algorithms.
IV-B Case Study 2: Autonomous Vehicles in Lane Merging
Figure 3: Trajectories of the lane-merging case study. (a) Under complete information, safe lane merging is ensured as the red car yields to the more aggressive blue car. (b) Under incomplete information, where each agent assumes the other is an expert during intent inference, a collision occurs. (c) N-PACE results show that when agents infer each other’s aggressiveness while also accounting for their peer’s learning, safe lane merging is achieved, with the red car yielding to the blue car.
[…]
$\dot{x}_{1}$ $=v_{1},\quad\dot{v}_{1}=a_{1},\quad\dot{x}_{2}=v_{2}\cos{\theta_{2}},\quad\dot{y}_{2}=v_{2}\sin{\theta_{2}},$
$\dot{\theta}_{2}$ $=\delta_{2},\quad\dot{v}_{2}=a_{2},\quad g_{blue}=0.1(x_{1}-25)^{2}+0.1v_{1}^{2}$
$+$ $\theta_{1}e^{-(d(t)-d_{\text{safe}})^{2}}+a_{1}^{2},\ g_{red}=1y_{2}^{2}+10\theta_{2}^{2}+0.1v_{2}^{2}$
$+$ $0.1(x_{2}-25)^{2}+\theta_{2}e^{-(d(t)-d_{\text{safe}})^{2}}+a_{2}^{2}+\delta_{2}^{2}.$
where $d(t)=(x_{1}-x_{2})^{2}+y_{2}^{2}$ is the distance between the two cars, $d_{\text{safe}}$ is the safe distance they aim to maintain, and $\theta_{k}$ is a parameter in the range $[0,100]$ that represents the aggressiveness of each agent (lower $\theta_{k}$ corresponds to more aggressive behavior).
Figure 4: Comparing parameter learning of each agent in case study 2 under N-PACE and ”treating others as Expert” with complete information, highlighting the superior performance of N-PACE and its fast convergence.
Although treating a learning agent as an expert and attempting to estimate their aggressiveness may not converge to the true parameter value, this approach does not always necessarily lead to a collision. In this example, we consider a specific scenario and initial conditions where the complete information game ensures safe lane merging, and our N-PACE algorithm also enables fast and accurate parameter learning; however, under the same initial conditions and learning dynamics, treating the peer agent as an expert leads to not only incorrect parameter learning but also collisions and safety issues. Each agent initially assumes that the other agent is a non-aggressive, safety-conscious driver with $\hat{\theta}^{k}_{0}=100$. However, both agents exhibit some level of aggressiveness: the first agent (the blue car) is highly aggressive with $\theta_{1}=5$, while the second agent (the red car) has a moderate aggressiveness level of $\theta_{2}=80$. As shown in Fig. 3, in the complete information game scenario, agent 2 (the red car) yields to agent 1 (the blue car), successfully avoiding a collision. However, in the incomplete information game, where agents treat their peers as experts, the biased learning process leads not only to incorrect parameter estimations but also to overlapping safety regions in their trajectories, resulting in safety risks. In contrast, N-PACE, due to its unbiased learning process, converges quickly to the true parameter values and avoids collisions. As a result, its behavior closely resembles that of the complete information game, with the red car yielding to the blue car, albeit with some latency.
IV-C Case Study 3: Autonomous Vehicles in an Interaction
Figure 5: Trajectory $(x,y)$ results of 300 simulations of the two-player incomplete information intersection game with random values of aggressiveness. Gray boxes indicate the collision area. (a) The complete information game results in no collisions. (b) The incomplete information game, where agents assume their peer is a complete information agent, results in 12 collision cases out of 300. (c) N-PACE results in only 4 collisions. (d) N-PACE combined with intent communication achieves 0 collisions and closest trajectories to those in the complete information game.
In our third case study, to extensively study N-PACE and the intent communication approach proposed in (7), we conduct Monte Carlo simulations of an intersection driving game between two agents under incomplete information, similar to the scenario shown in Fig. 1. In this case study, two autonomous vehicles attempt to navigate through an intersection while being unaware of each other’s aggressiveness parameter, $\theta_{k}$. The agents are only able to control their acceleration along with the direction of their movement $x,y$. The dynamic of the game and each agent cost function follow as:
$\dot{x}_{1}$ $=v_{1},\quad\dot{v}_{1}=a_{1},\quad\dot{y}_{2}=v_{2},\quad\dot{v}_{2}=a_{2},$
[…]
where $d(t)=x_{1}^{2}+y_{2}^{2}$ is the distance between two vehicles and $d_{safe}$ is the safe distance they try to maintain. The aggressiveness of each agent is also defined same as Case Study 2. The game continues for $5$ sec with a sampling time of $\Delta{T}=0.1$ sec, resulting in 50 samples.
In this example, each agent’s aggressiveness parameter, $\theta_{k}$, takes values in the range $[20,50]$, where a lower $\theta_{k}$ indicates a more aggressive agent. In our Monte Carlo study, we sampled 300 aggressiveness values uniformly from $[20,50]$ for each agent. Using these 300 sampled values, we ran the game 300 times under four different scenarios: 1) Complete information, 2) The expert peer agent assumption, 3) N-PACE, 4) N-PACE with intent communication, as proposed in Sec. III-E, for different intent communication weighting values $\eta=0.1,1,10$.
The trajectories from these 300 simulations for different scenarios are shown in Fig. 5, where we plot the $x$-position of the first agent against the $y$-position of the second agent. The gray box in the middle represents the collision area in the center of the intersection. As expected, under the complete information game scenario, no collisions were observed. However, in the scenario where each agent assumes its peer is a complete information/expert agent and attempts to infer their objective, we observed 12 collisions out of 300 simulations. In contrast, when agents used N-PACE to infer each other’s objectives, only 4 out of 300 trajectories resulted in collisions—a $66\%$ reduction compared to the baseline approach. It is important to note that having 4 collisions under N-PACE is not unexpected, as N-PACE also has a limited convergence speed during the interaction. This convergence speed depends on factors such as the chosen learning dynamics and the persistence of excitation in control signals. Interesting results were observed when we applied the proposed intent communication framework from Sec. III-E to this problem. In Fig. 5 (d), we present the trajectory results for $\eta=1$, where agents communicate their intent. Under this scenario, no collisions were observed in any of the 300 simulations. This was also the case for $\eta=10$, while only a single collision occurred for $\eta=0.1$. These results highlight the improvements achieved in both inference accuracy and safety assurance using the proposed intent communication framework built on top of N-PACE. Finally, Fig. 6 presents the log of the Mean Square learning Error (MSE) of both agents across all 300 Monte Carlo simulations. We use $\log(1+\text{MSE})$ for better visualization and comparison. The results clearly demonstrate the superiority of N-PACE over the approach where each agent treats the other as an expert, in terms of learning error. Additionally, they highlight how intent communication improves learning performance, particularly when the signaling is intensified using higher $\eta$.
Figure 6: Comparison of $\log(1+\text{MSE})$ learning error across 300 Monte Carlo simulations of the two-player incomplete information intersection game with random values of aggressiveness. The comparison includes various approaches, including N-PACE + Signaling with different values of the intent communication weighting factor $\eta$.
V CONCLUSIONS AND FUTURE WORK
In this paper, we presented an Nonlinear Peer-Aware Cost Estimation (N-PACE) framework for solving general-sum dynamic games, which relies on treating the peer agent as a learning agent while inferring its cost parameters. We demonstrated how this framework outperformed the approach of treating the peer agent as a complete-information agent in terms of task completion and safety assurance. Additionally, we proposed an intent communication framework built on top of N-PACE for signaling in incomplete information games.
N-PACE assumes access to the learning dynamics and initial estimates of the peer agent. While such information is often available in many multi-agent systems, there are scenarios—such as human-robot interactions—where the learning dynamics may be unknown. The question of ”learning the learning dynamics” presents an interesting research direction for future work, particularly in light of recent studies exploring the use of transformers for this purpose [10, 9]. Furthermore, the current framework is model-based, relying on iLQR solvers to solve the general-sum game. Integrating reinforcement learning-based controllers into N-PACE could open new research directions, enabling its application in model-free settings. Finally, while our numerical experiments demonstrate N-PACE’s reliable convergence, theoretical studies on its convergence in specific games and the impact of the intent communication weighting factor $\eta$ can further expand this work.